# FNOPE: Simulation-based inference on function spaces with Fourier Neural Operators

### v1

@@ Conclusion
We define a uniform distribution over the binary mask vectors with a fixed number of nonzero entries, $N_{\text{ds}}$. Suppose we are given a simulation $(\theta,l^{\theta},x,l^{x})$, where $\theta,l^{\theta}$ consist of $N_{\theta}$ points, and $x,l^{x}$ consist of $N_{x}$ points. We construct two random binary masks, $\mathbf{M}^{\theta}\in\{0,1\}^{N_{\theta}},\mathbf{M}^{x}\in\{0,1\}^{N_{x}}$, each with exactly $N_{\text{ds}}$ nonzero entries. We then remove the corresponding elements of $\theta,l^{\theta}$ where $\mathbf{M}_{\theta}$ is zero, and similarly remove the corresponding elements of $x,l^{x}$ where $\mathbf{M}_{x}$ is zero. For a minibatch, of simulations, we independently sample the masks $\mathbf{M}^{\theta}_{i},\mathbf{M}^{x}_{i}$ for each simulation. If $N_{\text{ds}}>N_{\theta}$ or $N_{\text{ds}}>N_{x}$, we leave the corresponding value and position vector unchanged. The value of $N_{\text{ds}}$ used in our work is reported for each experiment in Appendix S6.
@@
Another potential advantage of FNOPE (fix) is its computational efficiency. Consider the case of a parameter discretized uniformly in $D$ dimensions, with $L$ points per dimension, leading to a total of $L^{D}$ points. The computational cost of computing the spectral decomposition of the parameters using the FFT is $O(L^{D}\log L^{D})=O(DL^{D}\log L)$. Assuming the maximum number of modes in each dimension modeled by the FNO is $M$, this leads to $M^{D}$ total modes to compute. Therefore, the computational cost of the frequency-limited NUDFT is $O(M^{D}L^{D})$. For one-dimensional domains, the NUDFT may well be faster to compute than the FFT. However, for higher dimensions, the NUDFT scales exponentially with both $M$ and $L$. This discrepancy increases with both the dimensionality of the domain, and the number of modes $M$ modeled by the FNO blocks.
S3.4 Additional Parameters
The näive extension of the FMPE objective as stated in Sec. 3.3 is to minimize the $L^{2}$ loss $||v^{\phi}-u_{t}||^{2}$, where $v^{\phi}=[v^{\phi}_{l^{\theta}},v^{\phi}_{\eta}]$ and $u_{t}=[u_{t}(\xi_{t}|\theta),u_{t}(z_{t}|\eta)]$. However, the scale of the loss for the continuous parameters, $||v^{\phi}_{l^{\theta}}-u_{t}(\xi_{t}|\theta)||^{2}$ varies with the number of points in the discretization $l^{\theta}$, which we denote $N_{\theta}$. To ensure that the loss is balanced for the function- and vector-valued parameters, we in practice add their $L^{2}$ losses, normalized by their respective vector dimensionalities. That is, we minimize
$\mathbb{E}\frac{1}{N_{\theta}}||v^{\phi}_{l^{\theta}}-u_{t}(\xi_{t}|\theta)||^{2}+\frac{1}{N_{\eta}}||v^{\phi}_{\eta}-u_{t}(z_{t}|\eta)||^{2},$
where $N_{\eta}$ is the dimensionality of $\eta$. The expectation is over the same random variables as in the loss $\mathcal{L}_{2}$ in Eq. 4.
@@
For spectral NPE/FMPE we first apply a Fourier transformation to the parameters, take the first $M$ Fourier modes and expand these $M$ complex values to a real vector of dimension $2M$ representing the real and imaginary parts (for two dimensional data this results in $2M^{2}$ parameters). After inferring the posterior of the parameters in Fourier space and sampling from it, we apply the inverse Fourier transform to get samples in the spatial domain.
@@
For NPE (spectral) we infer the posterior over the coefficients of the first $M$ Fourier modes following the spectral preprocessing described above. We use NPE [3] with normalizing flows. We do not apply spectral preprocessing to the observations but pass raw observations through an embedding net as it is common practice in NPE.
@@
For the mass balance experiment (Sec. 4.5), we also compare to the approach of Moss et al. [39], which we refer to as NPE (raw). This approach infers the mass balance parameters on a fixed discretization of 50 gridpoints. The authors use a Neural Spline Flow [54] with 5 transformations, two residual blocks of 50 hidden units each, ReLU nonlinearities and 10 bins. The embedding net used to embed the 441-dimensional observation is a CNN with two convolutional layers of kernel size 5, with ReLU activations and max pooling of kernel size 2. The convolutional layers are followed by two linear layers with 50 hidden units and output dimension 50. The same settings are used in the 500-dimensional experiment (Appendix S7). Training is performed with a batch size of 200 and an Adam optimizer with learning rate of 0.0005 [55].
@@
As with NPE (spectral), in this approach we apply spectral preprocessing to the parameters, and infer the coefficients of the top $M$ Fourier modes, but process the observations directly using an embedding net. We use MLPs to estimate the flows, as in Wildberger et al. [9], as implemented in the sbi toolbox [29]. This implementation also uses the rectified flow [18] objective for FMPE, and we use independent Gaussian noise as the noise distribution. The flow networks are conditioned on time by concatenating the time to the inputs.
@@
For the SIRD experiment, we apply Simformer with the same settings as in Gloeckler et al. [30]. That is, we use a transformer model with a token dimension of $50$, 8 layers, and $4$ heads. The widening factor is 3, and the training was performed with a batch size of 1000 and an Adam optimizer. We train Simformer to learn all conditionals, and so uniformly draw between the posterior, joint, and likelihood masks, as well as two random masks drawn from Bernoulli distributions with $p=0.3$ and $p=0.7$ respectively. For both SIRD experiments, where we evaluate on 20 and 40 time points respective, we use the same Simformer model which is trained using 20 randomly sampled time points.
@@
The Gaussian Simulator is inspired by Lueckmann et al. [31], but instead of a 10 dimensional Gaussian distribution with independent dimensions we expanded the problem to 1000 dimensions and use a Gaussian Process prior (see below). Draws from the simulator are still drawn independently per dimension as $x\sim\mathcal{N}(\theta,\sigma^{2}\boldsymbol{I})$, where $\sigma^{2}=0.1$ (as in [31]).
@@
The prior is defined as a Gaussian process $\mathcal{GP}$ on $[0,1]$ with an equidistant discretization with 1000 timepoints. A draw from the prior is therefore defined as $\theta\sim\mathcal{GP}(0,k(\cdot,\cdot))$, where $k$ is the squared exponential kernel, $k(t,t^{\prime})=\exp(-\frac{(t-t^{\prime})^{2}}{2l^{2}})$. We set the lengthscale $l=0.05$ and variance to 1.
@@
Similar to Gloeckler et al. [30] we extend the SIRD (Susceptible, Infected, Recovered, Deceased) model to have a time-dependent contact rate. Compared to the classical SIR framework the model additionally incorporates a Deceased (D) population. Similar models were explored by Chen et al. [33], Schmidt et al. [34]. This addition is important for modeling diseases with significant mortality rates. The SIRD model, including a time-dependent contact rate $\beta(t)$, is defined by the following set of differential equations:
@@
We impose the same prior as in Gloeckler et al. [30]: the global variables $\gamma$ and $\mu$ are drawn from a Uniform distribution, $\gamma,\mu\sim\text{Unif}(0,0.5)$. For the time-dependent contact rate we define a Gaussian process prior which is further transformed by a sigmoid function to ensure that $\beta(t)\in[0,1]$ for all $t$. For the Gaussian process we use a RBF kernel $k$ defined as $k(t,t^{\prime})=\exp(-\frac{\lVert t-t^{\prime}\rVert^{2}}{2\cdot 7^{2}})$.
@@
MSE as well as SBC EoD is based on 100 observations with 1000 posterior samples each (Fig. 4c and Fig. S1c). To calculate the posterior predictive, we sample the initial condition $I(0)$ from the Simformer prediction for both methods, as opposed to the prior defined above, to match the setting of Gloeckler et al. [30].
@@
Details for the Darcy model are already given in the main text. We additionally scale the log-permebility $a$ by the scale factor of 1000 before taking the exponential - this results in permeabilities which produce sufficiently variable solutions using the Darcy flow simulator, and the permeabilities on the same scale as reported in Lim et al. [37]. The simulation output is additionally corrupted by an independent Gaussian observational noise per pixel $\zeta_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$. We set $\sigma_{i}=\mathbb{E}[u_{i}^{2}/30]$, where the expectation is per simulation batch, resulting in a signal to noise ratio (SNR) of 30.
@@
All metrics are calculated over a test set of 10 observations (Fig. 5 b-d). For MSE and SBC EoD, and 100 posterior samples were used for each observation and for each method. Finally, for SBC EoD, we use a subset of 50 pixels of the full $129\times 129$ as the marginals used for the reducing functions. The same pixels were used across all methods and all observations.
@@
We use a tracer transport [56, 57] simulator as described in Moss et al. [39]. This model takes in spatially varying surface accumulation rates $\dot{a}(l)$, which are related to the basal melt rates $\dot{b}(l)$ through the total balance condition $\dot{m}(l)=\dot{a}(l)-\dot{b}(l)$. $\dot{m}(l)$ is known and fixed across all simulations.
@@
where $u(l)$ is the known velocity profile of the ice shelf which is fixed across simulations. Additional layers are added at the top of the ice shelf and removed from the bottom according to $\dot{a}(l),\dot{b}(l)$ accordingly. The noise model approximates the observation noise of the radar measurement given an assumed density profile of the ice shelf. At the final timestep $T$, the layer that mostly closely matches the ground truth observation $x^{o}$ according to the $L^{2}$ norm is selected as the simulator output. Full details are described in Moss et al. [39].
@@
For SBC EoD, as well as the predictive MSE on synthetic test simulations, we use 100 test observations and sample 10 posterior samples for each observations, for each method. The real test data consists of one field observation (shown in Fig. 6a-b), and the posterior predictive was estimated using 1000 posterior samples.
@@
For all the baseline methods, we train the networks using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 50 modes, leading to 100 parameters to learn, and a pad width of 200 for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a Neural Spline Flow (NSF) with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 5 linear layers with 64 hidden dimensions to estimate the flow, with ELU activations. In both cases, we embedd the 1000 dimensional observation into a 40-dimensional vector using an MLP with 2 layers and 50 hidden units and RELU activations. For FMPE (raw), we use an MLP with 5 layers and 64 hidden features, with ELU activations.
@@
The overall number of trainable parameters for each network in this experiment is summarized in Table S1.
Table S1: Linear Gaussian Network sizes. The number of trainable parameters for each of the methods in the Linear Gaussian task.
Method ##\## Trainable Params
FNOPE 109,797
FNOPE (fix) 108,581
NPE (spectral) 566,590
FMPE (spectral) 34,212
FMPE (raw) 149,032
@@
For all the baseline methods, we train the network using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 16 modes, leading to $5122\times 16^{2}=$ 512 parameters to learn, and a pad with a width of 20 in each dimension for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a NSF with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 8 layers with 256 hidden dimensions to estimate the flow, with ELU activations. For FMPE (raw), we use an MLP with 8 layers and 256 hidden features, with ELU activations. All baseline methods embed the observation with a CNN embedding net into a 100-dimensional vector using 4 convolutional layers with kernel size 5 followed by max pooling of kernel size 2, followed by a 4-layer MLP with 100 hidden units, with RELU nonlinearities throughout.
@@
For all the baseline methods, we train the network using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 10 modes, leading to 20 parameters to learn, and a pad width of 20 for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a NSF with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 5 linear layers with 64 hidden dimensions to estimate the flow, with ELU activations. For FMPE (raw), we use an MLP with 5 layers and 64 hidden features, with ELU activations. For all baseline methods, we used the same embedding as Moss et al. [39], which was a CNN embedding the 441-dimensional observation into a 50-dimensional vector using 2 convolutional layers with kernel size 5 followed by max pooling of kernel size 2, followed by a 2-layer MLP with 50 hidden units, with RELU nonlinearities throughout. The configuration of NPE (raw) is described in Appendix S4.3.
For FNOPE  we use 10 Fourier modes for the FNO blocks. The network is made of 5 FNO blocks with 16 channels, while the context is embedded into 8 channels. We train for a maximum of 1000 epochs with an early patience of 50. We use a training batch size of 200 and a learning rate of 0.001. We do not include the data augmentation procedure for this experiment, as the discretizations of both observations and parameters was fixed to the setting of [39]. We still include positional embedding due to the parameter and observations being discretized differently to one another: the architecture included 4 channels for positional embedding, and 4 channels for time embedding. All nonlinearities are GELUs.

### v2

@@ Conclusion
We define a uniform distribution over the binary mask vectors with a fixed number of nonzero entries, $N_{\text{ds}}$. Suppose we are given a simulation $(\theta,l^{\theta},x,l^{x})$, where $\theta,l^{\theta}$ consist of $N_{\theta}$ points, and $x,l^{x}$ consist of $N_{x}$ points. We construct two random binary masks, $\mathbf{M}^{\theta}\in\{0,1\}^{N_{\theta}},\mathbf{M}^{x}\in\{0,1\}^{N_{x}}$, each with exactly $N_{\text{ds}}$ nonzero entries. We then remove the corresponding elements of $\theta,l^{\theta}$ where $\mathbf{M}_{\theta}$ is zero, and similarly remove the corresponding elements of $x,l^{x}$ where $\mathbf{M}_{x}$ is zero. For a minibatch of simulations, we independently sample the masks $\mathbf{M}^{\theta}_{i},\mathbf{M}^{x}_{i}$ for each simulation. If $N_{\text{ds}}>N_{\theta}$ or $N_{\text{ds}}>N_{x}$, we leave the corresponding value and position vector unchanged. The value of $N_{\text{ds}}$ used in our work is reported for each experiment in Appendix S6.
@@
Another potential advantage of FNOPE (fix) is its computational efficiency. Consider the case of a parameter discretized uniformly in a $D$-dimensional domain, with $L$ points per dimension, leading to a total of $L^{D}$ points. The computational cost of computing the spectral decomposition of the parameters using the FFT is $O(L^{D}\log L^{D})=O(DL^{D}\log L)$. Assuming the maximum number of modes in each dimension modeled by the FNO is $M$, this leads to $M^{D}$ total modes to compute. Therefore, the computational cost of the frequency-limited NUDFT is $O(M^{D}L^{D})$. For one-dimensional domains, the NUDFT may well be faster to compute than the FFT. However, for higher dimensions, the NUDFT scales exponentially with both $M$ and $L$. This discrepancy increases with both the dimensionality of the domain, and the number of modes $M$ modeled by the FNO blocks.
S3.5 Additional Parameters
The näive extension of the FMPE objective as stated in Sec. 3.3 is to minimize the $L^{2}$ loss $v^{\phi}-u_{t}||^{2}$, where $v^{\phi}=[v^{\phi}_{l^{\theta}},v^{\phi}_{\eta}]$ and $u_{t}=[u_{t}(\xi_{t}|\theta),u_{t}(z_{t}|\eta)]$. However, the scale of the loss for the continuous parameters, $||v^{\phi}_{l^{\theta}}-u_{t}(\xi_{t}|\theta)||^{2}$ varies with the number of points in the discretization $l^{\theta}$, which we denote $N_{\theta}$. To ensure that the loss is balanced for the function- and vector-valued parameters, we in practice add their $L^{2}$ losses, normalized by their respective vector dimensionalities. That is, we minimize
$\mathbb{E}\left[\frac{1}{N_{\theta}}||v^{\phi}_{l^{\theta}}-u_{t}(\xi_{t}|\theta)||^{2}+\frac{1}{N_{\eta}}||v^{\phi}_{\eta}-u_{t}(z_{t}|\eta)||^{2}\right],$
where $N_{\eta}$ is the fixed dimensionality of $\eta$. The expectation is over the same random variables as for the statement of the loss $\mathcal{L}_{2}$ in Sec. 3.3.
@@
For spectral NPE/FMPE we first apply a Fourier transformation to the parameters, take the first $M$ Fourier modes and expand these $M$ complex values to a real vector of dimension 2M12M-1 representing the real and imaginary parts (the imaginary part of the first component is always 0, hence it is discarded). For two dimensional data, the same preprocessing results in $2M^{2}-1$ parameters. After inferring the posterior of the parameters in Fourier space and sampling from it, we apply the inverse Fourier transform to get samples in the spatial domain.
@@
For NPE (spectral) we infer the posterior over the coefficients of the first $M$ Fourier modes following the spectral preprocessing described above. We use NPE [greenberg2019npec] with normalizing flows. We do not apply spectral preprocessing to the observations but pass raw observations through an embedding net as it is common practice in NPE.
@@
For the mass balance experiment (Sec. 4.5), we also compare to the approach of moss2023simulation, which we refer to as NPE (raw). This approach infers the mass balance parameters on a fixed discretization of 50 gridpoints. The authors use a Neural Spline Flow with 5 transformations, two residual blocks of 50 hidden units each, ReLU nonlinearities and 10 bins. The embedding net used to embed the 441-dimensional observation is a CNN with two convolutional layers of kernel size 5, with ReLU activations and max pooling of kernel size 2. The convolutional layers are followed by two linear layers with 50 hidden units and output dimension 50. The same settings are used in the 500-dimensional experiment (Appendix S7.4). Training is performed with a batch size of 200 and an Adam optimizer with learning rate of 0.0005.
@@
As with NPE (spectral), in this approach we apply spectral preprocessing to the parameters, and infer the coefficients of the top $M$ Fourier modes, but process the observations directly using an embedding net. We use MLPs to estimate the flows, as in wildberger2023fmpe, as implemented in the sbi toolbox [BoeltsDeistler_sbi_2025]. This implementation also uses the rectified flow [liu2023rectified] objective for FMPE, and we use independent Gaussian noise as the noise distribution. The flow networks are conditioned on time by concatenating the time to the inputs.
@@
For the SIRD experiment, we apply Simformer with the same settings as in glockler2024simformer. That is, we use a transformer model with a token dimension of $50$, 8 layers, and $4$ heads. The widening factor is 3, and the training was performed with a batch size of 1000 and an Adam optimizer. We train Simformer to learn all conditionals, and so uniformly draw between the posterior, joint, and likelihood masks, as well as two random masks drawn from Bernoulli distributions with $p=0.3$ and $p=0.7$ respectively. For both SIRD experiments, where we evaluate on 20 and 40 time points respective, we use the same Simformer model which is trained using 20 randomly sampled time points.
@@
The Gaussian Simulator is inspired by lueckmann2021benchmarking, but instead of a 10 dimensional Gaussian distribution with independent dimensions we expanded the problem to 1000 dimensions and use a Gaussian Process prior (see below). Draws from the simulator are still drawn independently per dimension as $x\sim\mathcal{N}(\theta,\sigma^{2}\boldsymbol{I})$, where $\sigma^{2}=0.1$ (as in [lueckmann2021benchmarking]).
@@
The prior is defined as a Gaussian process $\mathcal{GP}$ on $[0,1]$ with an equidistant discretization with 1000 timepoints. A draw from the prior is therefore defined as $\theta\sim\mathcal{GP}(0,k(\cdot,\cdot))$, where $k$ is the squared exponential kernel, $k(t,t^{\prime})=\exp(-\frac{(t-t^{\prime})^{2}}{2l^{2}})$. We set the lengthscale $l=0.05$ and variance to 1. Only in the ablation experiments (Fig. S1a,b) we changed the lenghtscale $l$ to 0.005.
@@
Similar to glockler2024simformer we extend the SIRD (Susceptible, Infected, Recovered, Deceased) model to have a time-dependent contact rate. Compared to the classical SIR framework the model additionally incorporates a deceased (D) population. Similar models were explored by chen2020timeSIR, schmidt2021probabilistic. This addition is important for modeling diseases with significant mortality rates. The SIRD model, including a time-dependent contact rate $\beta(t)$, is defined by the following set of differential equations:
@@
We impose the same prior as in glockler2024simformer: the global variables $\gamma$ and $\mu$ are drawn from a uniform distribution, $\gamma,\mu\sim\text{Unif}(0,0.5)$. For the time-dependent contact rate we define a Gaussian process prior which is further transformed by a sigmoid function to ensure that $\beta(t)\in[0,1]$ for all $t$. For the Gaussian process we use a RBF kernel $k$ defined as $k(t,t^{\prime})=\exp(-\frac{\lVert t-t^{\prime}\rVert^{2}}{2\cdot 7^{2}})$.
@@
MSE as well as SBC EoD is based on 100 observations with 1000 posterior samples each (Fig. 4c and Fig. S2c). To calculate the posterior predictive, we sample the initial condition $I(0)$ from the Simformer prediction for both methods, as opposed to the prior defined above, to match the setting of glockler2024simformer.
@@
Details for the Darcy model are already given in the main text. We additionally scale the log-permebility $a$ by the scale factor of 1000 before taking the exponential - this results in permeabilities which produce sufficiently variable solutions using the Darcy flow simulator, and the permeabilities on the same scale as reported in lim2025DDO. The simulation output is additionally corrupted by an independent Gaussian observational noise per pixel $\zeta_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$. We set $\sigma_{i}=\mathbb{E}[u_{i}^{2}/30]$, where the expectation is per simulation batch, resulting in a signal to noise ratio (SNR) of 30.
@@
All metrics are calculated over a test set of 10 observations (Fig. 5 b-d). For MSE and SBC EoD, 100 posterior samples were used for each observation and for each method. Finally, for SBC EoD, we use a subset of 50 pixels of the full $129\times 129$ as the marginals used for the reducing functions. The same pixels were used across all methods and all observations.
@@
We use the same simulator as described in moss2023simulation. This model takes in spatially varying surface accumulation rates $\dot{a}(l)$, which are related to the basal melt rates $\dot{b}(l)$ through the total balance condition $\dot{m}(l)=\dot{a}(l)-\dot{b}(l)$. $\dot{m}(l)$ is known and fixed across all simulations.
@@
where $u(l)$ is the known velocity profile of the ice shelf which is fixed across simulations. Additional layers are added at the top of the ice shelf and removed from the bottom according to $\dot{a}(l),\dot{b}(l)$ accordingly. The noise model approximates the observation noise of the radar measurement given an assumed density profile of the ice shelf. At the final timestep $T$, the layer that most closely matches the ground truth observation $x^{o}$ according to the $L^{2}$-norm is selected as the simulator output. Full details are described in moss2023simulation.
@@
For SBC EoD, as well as the predictive MSE on synthetic test simulations, we use 100 test observations and sample 10 posterior samples for each observations and for each method. The real test data consists of one field observation (shown in Fig. 6a-b), and the posterior predictive was estimated using 1000 posterior samples.
@@
For all the baseline methods, we train the networks using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 50 modes, leading to 100 parameters to learn, and a pad width of 20 for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a Neural Spline Flow (NSF) with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 5 linear layers with 64 hidden dimensions to estimate the flow, with ELU activations. In both cases, we embed the 1000 dimensional observation into a 40-dimensional vector using an MLP with 2 layers and 50 hidden units and RELU activations. For FMPE (raw), we use an MLP with 5 layers and 64 hidden features, with ELU activations.
@@
For all the baseline methods, we train the network using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 16 modes, leading to $2\times 16^{2}=512$ parameters to learn, and a pad width of 20 in each dimension for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a NSF with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 8 layers with 256 hidden dimensions to estimate the flow, with ELU activations. For FMPE (raw), we use an MLP with 8 layers and 256 hidden features, with ELU activations. All baseline methods embed the observation with a CNN embedding net into a 100-dimensional vector using 4 convolutional layers with kernel size 5 followed by max pooling of kernel size 2, followed by a 4-layer MLP with 100 hidden units, with RELU nonlinearities throughout.
@@
For all the baseline methods, we train the network using an Adam optimizer with a learning rate of 0.0001, and a batch size of 200. For NPE/FMPE (spectral), we use 10 modes, leading to 20 parameters to learn, and a pad width of 20 for the spectral preprocessing (Appendix S4.1). For NPE (spectral) the density estimator is a NSF with 2 residual blocks with 50 hidden dimensions each, 5 transforms, with RELU activations. For FMPE (spectral), we use an MLP with 5 linear layers with 64 hidden dimensions to estimate the flow, with ELU activations. For FMPE (raw), we use an MLP with 5 layers and 64 hidden features, with ELU activations. For all baseline methods, we used the same embedding as moss2023simulation, which was a CNN embedding the 441-dimensional observation into a 50-dimensional vector using 2 convolutional layers with kernel size 5 followed by max pooling of kernel size 2, followed by a 2-layer MLP with 50 hidden units, with RELU nonlinearities throughout. The configuration of NPE (raw) is described in Appendix S4.3.
For FNOPE  we use 10 Fourier modes for the FNO blocks. The network is made of 5 FNO blocks with 16 channels, while the context is embedded into 8 channels. We train for a maximum of 1000 epochs with an early patience of 50. We use a training batch size of 200 and a learning rate of 0.001. We do not include the data augmentation procedure for this experiment, as the discretizations of both observations and parameters was fixed to the setting of [moss2023simulation]. We still include positional embedding due to the parameter and observations being discretized differently to one another: the architecture included 4 channels for positional embedding, and 4 channels for time embedding. All nonlinearities are GELUs.
@@
Appendix S7 Ablation experiments
S7.1 Linear Gaussian
To investigate the influence of different hyperparameters on the performance of FNOPE, we ran several ablation experiments. First, we studied the performance of FNOPE in the deliberately insufficient setting where the prior distribution over the parameter contains higher-frequency modes than what is modeled by the FNO blocks. To this end, we changed the lengthscale of the prior distribution for the Linear Gaussian task (Sec. 4.2) by a factor of $10$ to $0.005$, resulting in higher frequency components for the parameter $\theta$. We then trained FNOPE with a varying number of spectral modes in the FNO blocks. With a lower number of modes, the performance of FNOPE degrades, and the posterior samples clearly miss the high frequency components present in the ground truth observations (Fig. S1a,b). Second, we varied the number of unmasked points in the FNOPE training procedure (Sec. 3.1). While the SWD decreases with the number of unmasked points, it saturates at 512 unmasked points, which is approximately half of the observed data (Fig. S1c). We then consider changing the lengthscale heuristic used to define the base distribution for FNOPE (Sec. 3.1), by defining $l=L_{0}/M$ for the number of modes $M$ and some lengthscale scaling factor $L_{0}$. In the extreme case $L_{0}=0$, this corresponds to the base distribution sampling uncorrelated white noise (WN). The lengthscale heuristic used in our work corresponds approximately to $L_{0}=4/\pi\approx 1.27$. In contrast to the other hyperparameter ablation experiments, varying the lengthscale scaling factor does not seem to impact the performance of FNOPE considerably for this task. FNOPE performs well over a wide range of lengthscale scaling factors (Fig. S1d). Finally, we consider a version of FNOPE which keeps the masking scheme as described in Sec. 3.1, but without adding positional noise to the remaining positions, which effectively only sees positions on the simulation grid during training (FNOPE (no jitter)). On this task, we evaluate on an equispaced grid, and we see that FNOPE (no jitter) performs similarly to FNOPE  (Fig. S1e). Therefore, the performance of FNOPE does not degrade from the inclusion of positional noise in this task.
Figure S1: Linear Gaussian ablation experiments. (a) Performance for $10^{4}$ training samples in terms of SWD for the Linear Gaussian experiments with varying number of Fourier modes in the FNO block. Note that we changed the lengthscale of the simulator prior by a factor of 10 to $0.005$ (Appendix S5. (b) Power analysis of FNOPE (fix) samples and ground truth $x$ for different number of used modes in the FNO block. (c) Performance in terms of SWD for different numbers of masked points in FNOPE. (d) Influence of the lengthscale of the noise process on the SWD. “WN” corresponds to white noise (an uncorrelated Gaussian distribution). (e) Same as Fig. 3 with a version of FNOPE in which we omit the adding of positional noise (FNOPE (no jitter)).
S7.2 SIRD
First, we compare the performance of FNOPE to Simformer on the SIRD task when using observations at 20 (instead of 40) randomly sampled times (Fig. S2). The performance of FNOPE slightly degrades, while Simformer still performs robustly. Second, we adapt the lengthscale heuristic used to define the base distribution for FNOPE (Sec. 3.1), by defining $l=L_{0}/M$ for the number of modes $M$ and some lengthscale scaling factor $L_{0}$. The lengthscale heuristic used in our work corresponds approximately to $L_{0}=4/\pi\approx 1.27$. We observe that FNOPE performs robustly for a wide range of lengthscale scaling factors.
Figure S3: SIRD ablation experiments. Influence of the lengthscale of FNOPE’s base distribution (Sec. 3.1) on posterior quality in terms of MSE of posterior predictive simulations and Simulation-based calibration error of diagonal (SBC EoD).
S7.3 Darcy Flow
In the Darcy task, we ablate the number of modes used in the FNO block for FNOPE. We see that while the MSE of posterior predictive simulation and posterior calibration as measured by the SBC EoD of marginal distributions is not strongly affected by the number of modes used, the posterior log-probability per pixel increases with the increased model capacity (Fig. S4a). This is expected, as more modes allow FNOPE  to estimate a more constrained posterior distribution, leading to higher posterior densities. Second, we adapt the lengthscale heuristic used to define the base distribution for FNOPE (Sec. 3.1), by defining $l=L_{0}/M$ for the number of modes $M$ and some lengthscale scaling factor $L_{0}$. The lengthscale heuristic used in our work corresponds approximately to $L_{0}=4/\pi\approx 1.27$. We see that increasing the lengthscale scaling factor improves the calibration of the posterior learned with FNOPE. We observe that FNOPE achieves its best performance in terms of posterior log-probability when the base distribution lengthscale scaling factor is set to a value around our lengthscale heuristic (Fig. S4b).
Figure S4: Darcy ablation experiments (a) Influence of number of used modes in the FNO block on different performance measures. The original experiments used 32 modes. Measures are the same as in Fig. 5b-d. (b) Influence of the noise length scale on different performance measures. The original experiments used a lengthscale of $\approx 1.27$. Measures are the same as in Fig. 5b-d.
S7.4 Mass balance rates of Antarctic Ice Shelves
We repeat the Antarctic ice mass balance experiment (Sec. 4.5), without downsampling the parameter space from simulation to inference, i.e. inferring the full 500-dimensional posterior distribution over the mass balance parameters. Overall, we observed that FNOPE maintains its performance on this higher-dimensional problem (Fig. S5). For low simulation budgets, the performance of FNOPE exceeds the other methods in terms of predictive MSE for both synthetic and real observations. As expected, spectral baseline methods significantly outperform the other baselines, especially at low simulation budgets.
Figure S5: Mass balance rates of ice shelves. Inference of mass balance rates where the parameter discretization uses 500 gridpoints (observation discretization remains unchanged). Performance measures on test simulations and the real observation, where NPE (raw) refers to the method used in moss2023simulation. Results for FMPE (raw) omitted as this baseline was not able to always produce samples within the prior bounds across all test observations. See Fig. 6 for the results based on 50 grid points.